# arrows.meta, chains, positioning

### v1

@@ I Introduction
Machine learning (ML) applications in healthcare drive towards automation of clinician tasks, enhance their capabilities and access to patient care [1]. The research community widely adopts advanced ML mechanisms such as Federated Learning due to their ability to maintain data privacy while building robust state-of-the-art server models [2]. Another ML mechanism Reinforcement Learning(RL) has shown great success in enhancing the prediction performance [3] and can assist in sequential decision-making tasks [4]. One of the most important applications of RL in dynamic environments is in the control of autonomous systems, such as robots and drones. These systems must be able to navigate and interact with their environment in real time, and RL can be used to train them to make optimal decisions based on sensor data. For example, RL can be used to train a robot to navigate through a cluttered environment, or a drone to fly in a safe and efficient manner.
In multiple domains, such as healthcare, traffic, and weather forecasting, early warning systems (EWS) are crucial to detect and diagnose issues early before they become more severe and challenging to handle. Real-time monitoring data is analyzed, and alerts are issued when certain thresholds are crossed, indicating a potential issue. Self-learning systems, like RL-based EWS, are particularly useful as they can adapt to changes in a condition over time and continually refine their predictions. By integrating an RL-based EWS into clinical decision-making, clinicians can make more informed decisions and take proactive measures to prevent adverse health events from occurring. RL-based EWS have been successful in predicting hospital readmissions, sepsis detection, and adverse drug events. In traffic, EWS can alert drivers to congestion, accidents, or other hazards, allowing them to plan alternative routes and avoid delays. In weather forecasting, EWS can provide advance notice of severe weather conditions, allowing individuals and communities to prepare and take necessary precautions to minimize damage and loss of life. Ultimately, EWS are vital in various domains to enable proactive measures, prevent damage, and enhance safety and quality of life.
RL environment with observation space defining states at each time step can be analogised to time-series data. Modelling time-series data for monitoring and predicting future states has been widely explored by deep learning models. Deep learning models, such as Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [5], play a significant role in predicting time-series data. These models are able to capture the temporal dependencies between time steps in the data, allowing them to make accurate predictions. The RNNs and LSTM architectures are designed to handle sequential data, making them well-suited for time-series prediction tasks.
Static temporal modelling, which is widely used in deep learning models like RNNs and LSTM networks, can help forecast and monitor potential issues and detect early warnings in various domains, including healthcare, traffic, and weather forecasting. These models can capture the temporal dependencies between time steps in the data and make accurate predictions. However, RNNs and LSTMs have some limitations when it comes to time-series prediction tasks, such as the assumption of equally spaced and ordered data and the lack of ability to incorporate graph structure where the data has a complex temporal structure.
This study proposes a novel framework called GraphRL, which combines reinforcement learning (RL) with graph-based neural networks to forecast the future states of a dynamic environment. Specifically, the framework uses a type of graph-based neural network called Temporal Graphical Convolutional Network (T-GCN) to capture the temporal dependencies between different time steps in a more natural way, by explicitly incorporating the graph structure of the data. One of the advantages of T-GCN is that it can also incorporate other forms of information, such as node and edge attributes, which can be useful for time-series prediction tasks where there is additional information associated with each time step. This is particularly important for deep learning challenges in Q-learning function approximation, where neural networks are used to approximate the Q-function. To evaluate the effectiveness of GraphRL, its performance is compared to other state-of-the-art models on a range of benchmark datasets. The results show that GraphRL outperforms the baseline models on most of the datasets, demonstrating the potential of using graph-based neural networks in RL. The contributions of this study are:
  • A novel and generic GraphRL framework with predictive and monitoring capabilities for early warnings in a complex environment.
  • A customised and generic RL environment to predict future states of a dynamic environment such as healthcare, traffic, and weather forecast.
  • A new paradigm of virtual monitoring of predicted states in RL, allowing for timely interventions by communicating with the appropriate team.
@@
The rest of the paper is organised as Section II presents the existing works on self-learning systems and prediction tasks. Section III presents the research problem. In Section IV, the proposed GraphRL framework and its algorithm are discussed. Section V presents the datasets used to evaluate the generic GraphRL framework and the baseline models are presented. Predictive RL environment and agent performances are presented and compared with the baseline models in Section VI-A. The proposed framework hyperparameters are fine-tuned with Bayesian optimisation in Section VII. Finally, the paper concludes in Section VIII.
@@ II-A Self Learning Systems
[6] proposed a framework representing a user as an RL agent and monitoring their behaviours in mobile health. The authors deployed two agents: a user agent to model the user decision-making process and an app agent to assist the user in achieving their goals by maximising the frequency of the user performing the behaviour. The proposed framework extracted different intervention strategies based on its self-learning experience. [7] adopted RL to model maladaptive eating behaviours for application-delivered mindfulness training intervention. [8] proposed Model Implants for Rapid Reflective Other-agent Reasoning/Learning (MIRROR) using RL to bootstrap human modelling. The MIRROR framework can quickly learn human behaviour from demonstrations. In personalising human behaviour, [9] discussed CalFit, a fitness app, built on the RL algorithm to generate personalised daily step goals by monitoring changes in behaviour and implementing dynamic goal setting and self-monitoring. Li et al. [10] propose a method that combines imitation learning with reinforcement learning to improve the success rate of autonomous driving in urban scenarios, which are more complex than simple scenarios such as highways and parking lots. The method decomposes the action space into low-level and high-level action spaces and uses a DQN-based reinforcement learning approach to execute specific imitation learning policies. The approach also introduces a new reward for high-level action selection.
@@ II-B Early Detection of Patient Deterioration
Early detection of vital signs deterioration is key to the timely invention and avoiding clinical deterioration in acutely ill patients in hospitals. Traditional patient monitoring is to report individual vital signs of patients which states their current clinical status. For example, vital signs such as temperature, pulse, respiratory rate, and mean arterial pressure are considered continuous predictors for emergency department patients [11]. New patient monitoring algorithms analyze multiple features from physiological signals. They can produce an index that serves as a predictive or prognostic measure for a specific critical health event or physiological instability [12, 13]. Ashrafian et al. presented a case series where wireless remote vital signs monitoring systems on medical and surgical wards could reduce the time to detect deteriorating patients [14]. Kellett et al. [15] presented numerous studies as evidence of the deterioration of patients is often missed even though it is usually preceded by worsening vital signs. Although vital signs are important to measure there remain, challenges as well as opportunities for new technology like wearable devices [15, 16]. Early recognition of patient deterioration in hospitals is crucial in reducing morbidity and preventing long-term disability. Early clinical prediction is typically based on manually calculated screening metrics that weigh these parameters, such as Early warning scores (EWS) [17]. EWS is considered one of the metrics to estimate early detection. There have been reports published that EWS are excellent predictors of cardiac arrest [18], ICU transfer [19], death in ICU [20] and as well as 30-day mortality and length of stay in ICU [21].
@@ II-C Vital Signs Prediction
[22] proposed two prediction models for mortality prediction and length of ICU stay respectively. The authors applied six traditional machine learning models to classify survived or not in mortality prediction, and ICU length of stay prediction with two machine learning models. Similarly, [23] proposed five predictive models in five different scenarios to predict vital signs such as heart rate, blood pressure, respiration rate and oxygen saturation of covid-19 patients in ICU. [24] proposed the patient’s clinical time-series variable prediction framework with multi-head transformers. The authors discussed four different scenarios: clinical time series forecasting, in-hospital mortality prediction, 24-hour mortality prediction and Length Of Stay prediction. [25] proposed a deep learning approach, DeepVS, for radio frequency-based vital signs sensing. The author combined 1D CNN and attention models to exploit local features and temporal correlations to estimate heart and respiration rates.
The use of self-learning algorithms in sequential decision-making tasks has garnered attention in the research community as they aid healthcare practitioners in making personalised interventions and prescribing appropriate medication. While deep learning methods such as CNN, RNN, and transformers have shown success in vital signs prediction, they have limitations such as assuming equally spaced and ordered data, and not being able to incorporate graph structure. Therefore, the aim of this study is to introduce an adaptive self-learning prediction framework that utilises the power of GNNs [26, 27] to capture temporal dependencies between timesteps and enable real-time monitoring with reinforcement learning.
@@ III Research Problem
The research problem is to overcome the deep learning challenges and predict future states of a complex and dynamic RL environment and learn the latent behaviour patterns of data adaptively.
Definition 1
Let $V$ be set of $n$ vital parameters $V_{t}=\{v^{1},v^{2},...,v^{n}\}$ at time $t$ determining the health status of a subject $S$. The vital parameters are continuous time-series data changing dynamically and inferring the subject’s health status. For time-series forecasting, the continuous vital parameters are split into time windows such as $T=\{V_{t-2},V_{t-1},V_{t},V_{t+1},V_{t+2},.....,V_{t+n}\}$. Non-linear models are trained with the historic data in time windows $\{V_{t-2},V_{t-1},V_{t}\}$ to predict future $\{V_{t+1},V_{t+2},.....,V_{t+n}\}$.
Definition 2
Based on the training, the subject $S$ has a set of learning agents following the Markov Decision Process (MDP) to monitor and learn the patterns of their vital parameters $V_{t}=\{v^{1},v^{2},...,v^{n}\}$ continuously. This can be defined as a 5-tuple $M$= ($S,A,P,R,\gamma$) constituting $S$ as finite state space, in which $s_{t}\epsilon S$ denotes the state of an agent at time $t$, $A$ denotes set of actions defined for the agent in which $a_{t}\epsilon A$ denotes the action taken by the agent at time $t$, $P$ is a Markovian transition function which denotes agent transits from state $s$ to state $s^{\prime}$ while performing an action $a$ through $P(s,a,s^{\prime}):S\times A\times S$. The transition function $P$ leads to a reward function $R:S\times A\rightarrow\Re$ which returns an immediate reward $R(s,a)$ for the action $a$ taken in a state $s$ defined in Equation 1.${\gamma}$ is a discount factor that focuses on immediate rewards instead of future rewards, which remains between 0 and 1.
${\displaystyle R(s_{t},a_{t})=\sum_{t=0}^{\infty}\gamma^{t}r_{t},}$ (1)
returns an immediate reward R(s, a) for the action taken in a state s defined in Equation 1
@@ IV-A1 T-GCN Forecast
Forecasting the future states of a subject before a few time steps can revolutionise the most dynamic industries such as gaming, healthcare, and so on by identifying the deteriorating state of the subject in the environment. To predict the future state in a reinforcement learning environment, a temporal graph convolutional network(T-GCN) is adopted. The graphical network is trained with past and current states at their timestamps in a supervised approach as shown in Equation 3. The training process also includes the features leading to those states.
@@ IV-A2 Static Spatial-temporal modelling
A two-layered graphical network is adopted for Spatial-temporal modelling, a spatial modelling layer is based on a graphical convolutional network (GCN), and a temporal layer based on recurrent neural networks (RNN) is configured. The spatial layer is responsible to capture spatial features among nodes which are input states $s_{t}^{i}\epsilon S$. This can be achieved by constructing Fourier transform filter and it acts on the graph nodes and its first-order neighbourhood. In this study, a static graph temporal signal is adopted in which the node positions in the graph remain the same and the label information is dynamic. The spatial layer is to set the static graph with nodes as input states $s_{t}^{i}\epsilon S$. The two-layered GCN model is defined in Equation 4.
@@
Temporal modelling is based on the RNN variant gated recurrent unit (GRU) [28] which has a simple structure and faster training ability. In the GRU model, an update gate $z_{t}$ controls the degree of information retrieved from the previous state and a reset gate $r_{t}$ controls the degree of ignoring the status information at the previous moment are configured as shown in Equation 5.
@@
In the training process, the T-GCN model predicts the future states at time $t+1,t+2,t+3,...,t+n$ as $\hat{Y_{t}}$ and compares with the real-data ${Y_{t}}$. This determines the loss function of the graph network [29] as shown in Equation 6. To avoid over-fitting problems in the training process, the loss function is optimised with L2 regularisation $L_{reg}$ and a hyperparameter $\lambda$.
@@ IV-B Algorithm I
1:Input: time series data $\mathcal{D}=\{s_{t-n},..,s_{t-2},s_{t-1},s_{t},\}$; a set of labels $\mathcal{K}=\{1,2,\dots,K\}$
2:Output: Predicted time series data of $\mathcal{K}$, a set of labels, in the form of states $\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}$.
@@
4:$Train(forecast\_model)\leftarrow forecast\_model(D)$
@@
6:$Initialization:\ observation\_space={s_{t}^{i}\epsilon S},action\_space={a_{t}\epsilon A},rewardR$
@@
The proposed GraphRL environment is defined in Algorithm 1, which configures the observation space, action space, and reward policy based on predicted states. The T-GCN model is defined in lines 1-3 to train and predict time series forecasting data. The class is initialized and boundaries for observation space, action space, rewards, and monitoring length are set in lines 4-5. Lines 6-10 describe the reward policy for learning agent actions and how agents are rewarded positively or negatively. Finally, lines 11-19 address monitoring length, visualization of agent performance, and resetting the environment if necessary.

### v2

# Graph-enabled Reinforcement Learning for Time Series Forecasting with Adaptive Intelligence

@@ I Introduction
The emergence of Machine Learning (ML) in healthcare signifies a paradigm shift towards automating clinician tasks and augmenting patient care capabilities [1]. Amidst the evolving ML landscape, Federated Learning has gained traction for preserving data privacy while constructing sophisticated server models [2]. Reinforcement Learning (RL), another ML strategy, has demonstrated substantial improvements in prediction performance and decision-making tasks [3, 4]. RL’s application is particularly noteworthy in controlling autonomous systems, such as robots and drones, training them to make optimal decisions in real-time based on environmental sensor data.
In various sectors, including healthcare, traffic, and weather forecasting, Early Warning Systems (EWS) play a pivotal role. They analyze real-time monitoring data and issue alerts for potential issues, facilitating proactive responses. RL-based EWS can adapt over time, refining their predictions and supporting clinical decision-making. This adaptability has proven effective in applications like predicting hospital readmissions and sepsis detection.
Figure 1: Graphical Abstract
Time-series data modeling, vital in monitoring and predicting future states, has seen advancements with deep learning models like Recurrent Neural Networks (RNNs) and Long Short-Term Memory (LSTM) networks [5]. These models are adept at capturing temporal dependencies, yet face limitations in handling irregularly structured data and integrating complex graph structures. This study introduces the GraphRL framework, an innovative amalgamation of RL and Temporal Graphical Convolutional Networks (T-GCN), aiming to surpass the constraints of traditional deep learning models in time-series prediction. GraphRL’s design facilitates handling complex temporal structures and incorporates additional information such as node and edge attributes, as depicted in Fig. 1. The GraphRL framework’s core contributions include:
  • A versatile framework capable of providing early warnings and monitoring in complex settings.
  • A customizable RL environment designed for effective forecasting in dynamic domains like healthcare and traffic systems.
  • A novel approach to virtual monitoring of predicted states in RL, enhancing decision-making and intervention capabilities.
Our comparative analysis with state-of-the-art models across various datasets showcases GraphRL’s superior performance, underscoring its potential as a versatile solution for time-series prediction challenges.
@@
The paper is organized as follows: Section II reviews existing literature on self-learning systems and prediction tasks. Section III outlines the research problem. The proposed GraphRL framework and its algorithm are detailed in Section IV. Section V describes the datasets and baseline models used for evaluation. The performance of the predictive RL environment and agent is compared with baseline models in Section VI. Section VII discusses the fine-tuning of the framework’s hyperparameters using Bayesian optimization. The paper concludes in Section VIII.
@@ II-A Self Learning Systems
Self-learning systems, particularly those utilizing Reinforcement Learning (RL), have seen significant advancements in various applications. For instance, Shin et al. [6] introduced a dual-agent framework in mobile health, effectively demonstrating user modeling and behavior intervention strategies. This work underscores the potential of RL in personalizing user experiences, a concept that aligns with our GraphRL framework’s goal of adaptive learning in dynamic environments. Similarly, Taylor et al. [7] applied RL in modeling maladaptive eating behaviors, further showcasing RL’s versatility in behavior prediction and modification. Chen et al. [8] developed the MIRROR framework, emphasizing the rapid learning capabilities of RL in human behavior modeling. These advancements set a precedent for our work in complex sequential decision-making tasks. Zhou et al.’s [9] CalFit app and Li et al.’s [10] method in autonomous driving highlight RL’s efficacy in personalized goal setting and complex urban scenario navigation, respectively, which are foundational to our GraphRL framework’s approach in handling dynamic and intricate patterns in data.
@@ II-B Early Detection of Patient Deterioration
In the healthcare domain, early detection of patient deterioration is vital. Traditional vital signs monitoring, as discussed by Asiimwe et al. [11] and evaluated by Scully et al. [12], Baig et al. [13], and others, has laid the groundwork for our study. These works highlight the importance of continuous monitoring and early warning systems (EWS), which are integral to GraphRL’s objective. The limitations in existing methods, such as the need for manual calculations and the inability to handle large, unstructured data effectively, are addressed in our framework through the integration of GNNs, which can process complex temporal data more efficiently.
@@ II-C Vital Signs Prediction
Vital signs prediction has been explored through various machine learning models. Alghatani et al. [14] and Youssef et al. [15] demonstrated the use of traditional machine learning in mortality prediction and vital signs forecasting, respectively. Harerimana et al.’s [16] work with multi-head transformers and Xie et al.’s [17] DeepVS model highlight the potential of deep learning in this domain. However, these methods often assume equally spaced and ordered data and lack the ability to incorporate complex graph structures, limitations our GraphRL framework aims to overcome.
II-D Temporal Graphical Convolutional Networks (T-GCN)
The integration of T-GCN within our GraphRL framework is pivotal. T-GCNs, known for their ability to capture temporal dependencies and complex relationships in graph-structured data, offer significant enhancements in processing time-series data [18]. This technology addresses limitations in traditional deep learning models by effectively managing irregular time intervals and integrating additional contextual information (such as node and edge attributes) for richer data representation [19]. The inclusion of T-GCN in GraphRL allows for a more nuanced understanding and prediction of dynamic systems [20], making it highly suitable for applications in healthcare monitoring, traffic forecasting, and weather prediction. The capability of T-GCNs to handle non-linear and complex temporal patterns [21] aligns with the core objectives of GraphRL, pushing the boundaries of current self-learning systems in real-world scenarios.
In summary, while existing works in self-learning systems, patient deterioration detection, and vital signs prediction have laid a strong foundation, our GraphRL framework aims to address their limitations by introducing a novel approach that combines the strengths of GNNs and RL. This approach allows for a more sophisticated handling of temporal dependencies and real-time monitoring, which is crucial in dynamic environments such as healthcare, traffic management, and weather forecasting.
The motivation behind the use of RL in our framework primarily arises from the need to tackle the challenges of multi-step time series prediction, where traditional supervised learning approaches may encounter limitations. Although supervised learning methods like GNN+Bert and GNN+TCN are indeed common and effective for time series forecasting, RL offers a unique advantage in dealing with situations where errors can accumulate over time, especially in dynamic environments. RL enables our predictive GraphRL Environment not only to forecast future states but also to actively influence decision-making, a capability particularly valuable in applications such as healthcare monitoring and the gaming industry.
@@ III Research Problem
The research problem addresses deep learning challenges in predicting future states of a complex and dynamic Reinforcement Learning (RL) environment and adaptively learning latent behavior patterns of data.
Definition 1 (Vital Parameters and Time-Series Forecasting)
In the context of our framework, we consider a set $V$ of $n$ vital parameters, denoted as $V_{t}=\{v^{1},v^{2},\ldots,v^{n}\}$, which represent continuous time-series data reflecting the health status of a subject $S$. These vital parameters are dynamic and change over time, providing valuable insights into the subject’s well-being. To facilitate time-series forecasting, we segment these continuous vital parameters into time windows, denoted as $T$, which encompasses data points from the past ($V_{t-2},V_{t-1},V_{t}$) and extends into the future ($V_{t+1},V_{t+2},\ldots,V_{t+n}$). Non-linear models are trained on historical data within the time windows $\{V_{t-2},V_{t-1},V_{t}\}$ to predict future values $\{V_{t+1},V_{t+2},\ldots,V_{t+n}\}$.
Definition 2 (Learning Agents and Markov Decision Process)
Following the training phase, subject $S$ is associated with a group of learning agents that operate based on the principles of the Markov Decision Process (MDP). This MDP is a 5-tuple denoted as $M=(S,A,P,R,\gamma)$, and it forms the foundation for continuous monitoring and pattern learning of the vital parameters $V_{t}=\{v^{1},v^{2},\ldots,v^{n}\}$. Here’s a breakdown of each component:
  • $S$ represents a finite state space, where $s_{t}\in S$ signifies the state of an agent at a specific time $t$,
  • $A$ is the set of actions available to each agent, and $a_{t}\in A$ represents the action taken by the agent at time $t$,
  • $P$ is a Markovian transition function $P(s,a,s^{\prime})$ that quantifies the probability of the agent transitioning from state $s$ to state $s^{\prime}$ while executing action $a$,
  • $R$ is a reward function $R:S\times A\rightarrow\mathbb{R}$ that provides an immediate reward $R(s,a)$ for the action $a$ performed in state $s$,
  • $\gamma$ is a discount factor, ranging between 0 and 1, which emphasizes immediate rewards over future rewards.
$R(s_{t},a_{t})=\sum_{t=0}^{\infty}\gamma^{t}r_{t},$ (1)
This equation returns the immediate reward $R(s,a)$ for the action taken in state $s$, as defined in Eq. 1.
@@ IV-A1 T-GCN Forecast
Forecasting the future states of a subject before a few time steps can revolutionise the most dynamic industries such as gaming, healthcare, and so on by identifying the deteriorating state of the subject in the environment. To predict the future state in a reinforcement learning environment, a temporal graph convolutional network(T-GCN) is adopted. The graphical network is trained with past and current states at their timestamps in a supervised approach as shown in Eq. 3. The training process also includes the features leading to those states.
@@
Eq. 2 describes a basic neural network neuron, with $y$ as the output, $f$ as the activation function, $b$ as the bias, $v_{i}$ as the input features, and $w_{i}$ as the weights; while Eq. 3 involves an ’Activation1’ function, computing a weighted sum of inputs and normalizing these outputs, possibly into a probability distribution, akin to a softmax function.
@@ IV-A2 Static Spatial-temporal modelling
A two-layered graphical network is adopted for Spatial-temporal modelling, a spatial modelling layer is based on a graphical convolutional network (GCN), and a temporal layer based on recurrent neural networks (RNN) is configured. The spatial layer is responsible to capture spatial features among nodes which are input states $s_{t}^{i}\epsilon S$. This can be achieved by constructing Fourier transform filter and it acts on the graph nodes and its first-order neighbourhood. In this study, a static graph temporal signal is adopted in which the node positions in the graph remain the same and the label information is dynamic. The spatial layer is to set the static graph with nodes as input states $s_{t}^{i}\epsilon S$. The two-layered GCN model is defined in Eq. 4.
@@
Temporal modelling is based on the RNN variant gated recurrent unit (GRU) [22] which has a simple structure and faster training ability. In the GRU model, an update gate $z_{t}$ controls the degree of information retrieved from the previous state and a reset gate $r_{t}$ controls the degree of ignoring the status information at the previous moment are configured as shown in Eq. 5.
@@
In the training process, the T-GCN model predicts the future states at time $t+1,t+2,t+3,...,t+n$ as $\hat{Y_{t}}$ and compares with the real-data ${Y_{t}}$. This determines the loss function of the graph network [23] as shown in Eq. 6. To avoid over-fitting problems in the training process, the loss function is optimised with L2 regularisation $L_{reg}$ and a hyperparameter $\lambda$.
@@ IV-B Predictive GraphRL Environment Algorithm
1:Input: time series data $\mathcal{D}=\{s_{t-n},\ldots,s_{t-2},s_{t-1},s_{t}\}$; a set of labels $\mathcal{K}=\{1,2,\dots,K\}$
2:Output: Predicted time series data of $\mathcal{K}$, a set of labels, in the form of states $\{s_{t+1},s_{t+2},s_{t+3},s_{t+4}\}$
@@
4:$Train(forecast\_model)\leftarrow forecast\_model(\mathcal{D})$
@@
6:$Initialization:observation\_space=\{s_{t}^{i}\in S\},action\_space=\{a_{t}\in A\},rewardR$
@@
Algorithm 1 outlines the creation of the Predictive GraphRL Environment, a crucial component of the proposed GraphRL framework. This environment leverages the T-GCN, chosen for its effectiveness in capturing spatial-temporal dynamics essential for complex systems. The T-GCN’s ability to forecast future states, in addition to analyzing current and historical data, makes it invaluable for critical applications like health monitoring and traffic management, where early detection and timely response are key. A significant feature of this algorithm is its reward mechanism, which is instrumental in guiding the learning process. By awarding rewards based on the suitability of the agent’s actions, the environment ensures that the agent’s policy is aligned with the primary objectives of accurate forecasting and effective intervention. This design was motivated by the need for a proactive system, capable of not only forecasting but also informing real-time decision-making. The algorithm is meticulously structured to set up the observation space, action space, and reward policy based on predicted states. The initial lines (1-3) justify the use of the T-GCN model, especially for its applicability in dynamic and nonlinear data contexts like vital sign monitoring. The subsequent lines (4-5) are dedicated to initializing the environment, forming the basis for RL-driven decision-making. The reward policy, detailed in lines 6-10, aligns with standard RL practices, promoting actions that yield beneficial outcomes. Finally, lines 11-19 focus on continuous monitoring and adaptation, a critical aspect for applications that demand real-time responsiveness, such as in healthcare scenarios. This algorithm represents a significant step in advancing the field of RL, moving from passive observation to an active role in shaping decisions.